# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v13

@@ VI Conclusion
In this paper, we introduce a novel gradient variant, referred to as the quadratic gradient, and implement a quadratic-gradient-enhanced version of Nesterov’s Accelerated Gradient (NAG) to train a logistic regression model in the encrypted domain. The proposed quadratic gradient is preferably constructed directly from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques. This approach exhibits strong potential for accelerating a broad range of optimization algorithms. In particular, it offers a viable alternative to traditional line-search procedures, can be incorporated into adaptive methods such as enhanced AdaGrad, and may serve to accelerate Newton’s method via embedded gradient descent steps.
As future work, we aim to establish a more rigorous theoretical analysis of the convergence properties of the enhanced algorithm. A more practical and application-driven direction involves systematically evaluating the performance of the enhanced method under different learning rate configurations.

### v14

@@ VI Conclusion
In this paper, we proposed a novel gradient variant, termed the quadratic gradient, and developed a quadratic-gradient-enhanced NAG for training logistic regression models in the encrypted domain.
The proposed quadratic gradient could be directlydirectly constructed from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques. This approach exhibits strong potential for accelerating a broad range of optimization algorithms. In particular, it offers a viable alternative to traditional line-search procedures and constitutes a game-changing approach to accelerating Newton’s method via embedded gradient-descent steps.
Open directions for future work include conducting a rigorous theoretical analysis of the enhanced algorithm’s convergence behavior and performing an application-oriented evaluation across diverse learning rate configurations.